The reaction goes to completion, but in the process of recovering the product, some of it was lost. The the percent yield for the reaction is 82.2%.

82.2%. How many grams of product are recovered?

How many grams of the excess reactant remain? Assume the reaction goes to completion.


1
Expert's answer
2022-03-19T02:45:25-0400

2KClO3→2KCl+3O2

From the reaction,it can be seen that;

2mol of KClO3= 3mol of O2

1 mol of KClO3=3/2 mol O2

=1.5 mol O2

The yield of the reaction is given in % i.e 85.5%

1 mil KClO3 generates 1.5 mol O2

Since the yield is 85.5%,the number of moles of O2 generated per mole of KClO3 is

1.5mol×85.8/100=1.2825mol

Number of moles ofO2=42.0g/32.0g

=1.31 mol.

Number of moles of KClO3

=1.31mol O2/1.2825 mol O2/KClO3

=1.68 mol KClO3

Weight of KClO3=1.35 mol KClO3×molecular weight of KClO3

=1.68×112.6g

=206g